Qd = 500 – 2P and Q5 = -100 + 3P

a. market equilibrium price is $_______________

b. market equilibrium quantity is______________

c. with government interference the current price is $80, the quantity supplied is___________ and the quantity demanded is___________________

d. how would you describe the market situation in c above and what would you expect to happen in this market?
Answer #1

Anewer - . Qd = 500-21 Os = -100 + 3P At the equilibrium, Quantity demanded - Quantity seepplied 500 - 21 =-100 +3P 2 600 =5P